\begin{thebibliography}{99}

\bibitem{putten1}
van Putten M H P M 1999 
Superradiance in a torus magnetosphere around a black hole 
\emph{Science} {\bf 284} 115

\bibitem{aguirre1}
Aguirre A N 2000
On the superradiance of spin-1 waves in an equatorial wedge around a Kerr hole 
\emph{ApJ} {\bf 529} L9

\bibitem{penrose1}
Penrose R 1969
\emph{Nuovo Cimento} {\bf 1} 252

\bibitem{bardeen1}
Bardeen J M, Press W H and Teukolsky S A 
Rotating black holes: Locally nonrotating frames, energy extraction, 
and scalar synchrotron radiation 
\emph{ApJ} {\bf 178} 347

\bibitem{wilkins1}
Wilkins D C 1972 
Bound geodesics in the Kerr metric 
\emph{Phys. Rev. D} {\bf 5} 814

\bibitem{wald1}
Wald R M 1974 
Energy limit on the Penrose process 
\emph{ApJ} {\bf 191} 231

\bibitem{christodoulou1}
Christodoulou D 1970 
Reversible and Irreversible transformations in black hole physics 
\emph{Phys. Rev. Lett.} {\bf 25} 1596

\bibitem{christodoulou2}
Christodoulou D and Ruffini R 1970
Reversible transformation of a charged black hole 
\emph{Phys. Rev. D} {\bf 4} 3552

\bibitem{teukolsky0}
Teukolsky S A 1972 
Rotating black holes: Separable wave equations for gravitational and 
electromagnetic perturbations 
\emph{Phys. Rev. Lett.} {\bf 29} 1114

\bibitem{teukolsky1}
Teukolsky S A 1973 
Perturbations of a rotating black hole. I. Fundamental equations for 
gravitational, electromagnetic, and neutrino-field perturbations 
\emph{ApJ} {\bf 185} 635

\bibitem{teukolsky2}
Press W H and Teukolsky S A 1973 
Perturbations of a rotating black hole. II. Dynamical stability of 
the Kerr metric
\emph{ApJ} {\bf 185} 649

\bibitem{teukolsky3}
Teukolsky S A and Press W H 1974 
Perturbations of a rotating black hole. III. Interaction of the hole 
with gravitational and electromagnetic radiation
\emph{ApJ} {\bf 193} 443

\bibitem{chandra0}
Chandrasekhar S 1983 
{\em The Mathematical Theory of Black Holes} 
(Oxford: Claredon Press)

\bibitem{aliev1}
Aliev A N and \"{O}zdemir N 2002 
Motion of charged particles around a rotating black hole in a magnetic field 
\emph{MNRAS} {\bf 336} 241

\bibitem{komissarov1}
Komissarov S S 2005 
Observations of the Blandford-Znajek and the MHD Penrose processes in 
computer simulations of black hole magnetospheres 
\emph{MNRAS} {\bf 359} 801

\bibitem{koide1}
Koide S, Shibata K, Kudoh T and Meier D L 2002 
Extraction of black hole rotational energy by a magnetic field and 
the formation of relativistic jets 
\emph{Science} {\bf 295} 1688

\bibitem{blandford1}
Blandford R D and Znajek R L 1977 
Electromagnetic extraction of energy from Kerr black holes 
\emph{MNRAS} {\bf 179} 433

\bibitem{znajek1}
Znajek R L 1977
Black hole electrodynamics and the Carter tetrad 
\emph{MNRAS} {\bf 179} 457

\bibitem{blandford2}
Blandford R D and Payne D G 1982 
Hydromagnetic flows from accretion discs and the production of radio jets 
\emph{MNRAS} {\bf 199} 883

\bibitem{znajek2}
Znajek R L 1978 
Charged current loops around Kerr holes 
\emph{MNRAS} {\bf 182} 639

\bibitem{znajek3}
Znajek R L 1978 
The electric and magnetic conductivity of a Kerr hole 
\emph{MNRAS} {\bf 185} 833

\bibitem{wilson1}
Wilson A S and Colbert E J M 1995  
The difference between radio-loud and radio-quiet active galaxies 
\emph{ApJ} {\bf 438} 62

\bibitem{xinli1}
Li L-X 2000
Electromagnetic energy for a charged Kerr black hole in 
a uniform magnetic field 
\emph{Phys. Rev. D} {\bf 61} 084033

\bibitem{wald2}
Wald R M 1974 
Black hole in a uniform magnetic field 
\emph{Phys. Rev. D} {\bf 10} 1680

\bibitem{king1}
King A R, Lasota J P and Kundt W 1975 
\emph{Phys. Rev. D} {\bf 12} 3037 

\bibitem{petterson1}
Petterson J A 1974 
Magnetic field of a current loop around a Schwarszchild black hole 
\emph{Phys. Rev. D} {\bf 10} 3166

\bibitem{petterson2}
Petterson J A 1975 
Stationary axisymmetric electromagnetic fields around a rotating black hole 
\emph{Phys. Rev. D} {\bf 12} 2218

\bibitem{chitre1}
Chitre D M and Vishveshwara 1975 
Electromagnetic field of a current loop around a Kerr black hole 
\emph{Phys. Rev. D} {\bf 12} 1538

\bibitem{bhmag1}
Ernst F J 1976 
Black holes in a magnetic universe 
\emph{J. Math. Phys.} {\bf 17} 54

\bibitem{bhmag1a}
Ernst F J and Wild W J 1976
Kerr black holes in a magnetic universe
\emph{J. Math. Phys.} {\bf 17} 182

\bibitem{bhmag2}
Hiscock W A 1981 
On black holes in magnetic universes 
\emph{J. Math. Phys.} {\bf 22} 1828

\bibitem{bhmag3}
D\'{i}az A G 1985 
Magnetic generalization of the magnetized Kerr-Newman black hole 
\emph{J. Math. Phys.} {\bf 26} 155

\bibitem{bhmag4}
Karas V and Vokrouchlick\'{y} 1991 
On interpretation of the magnetized Kerr-Newman black hole 
\emph{J. Math. Phys.} {\bf 32} 714

\bibitem{narayan1}
Narayan R and Quataert E 2005 
Black hole accretion 
\emph{Science} {\bf 307} 77 

\bibitem{semenov1}
Semenov V, Dyadechkin S and Punsly B 2004 
Simulations of jets driven by black hole rotation 
\emph{Science} {\bf 305} 978

\bibitem{wilms1}
Wilms J, Reynolds C S, Begelman M C, Reeves J, Molendi S, Staubert R and  
Kendziorra E 2001 
XMM-EPIC observation of MCG--6-30-15: Direct evidence for the extraction of 
energy from a spinning black hole ?
\emph{MNRAS} {\bf 328} L27

\bibitem{iwasawa1}
Iwasawa K, Fabian A C, Reynolds C S, Nandra K, Otani C, Inoue H, 
Hayashida K, Brandt W N, Dotani T, Kunieda H, Matsuoka M and Tanaka Y 1996 
The variable iron K emission line in MCG--6-30-15 
\emph{MNRAS} {\bf 282} 1038 

\bibitem{dabrowski1}
Dabrowski Y, Fabian A C, Iwasawa K, Lasenby A N and Reynolds C S 1997 
The profile and equivalent width of the X-ray iron emission line from 
a disc around a Kerr black hole 
\emph{MNRAS} {\bf 288} L11

\bibitem{shibata1}
Shibata M, Taniguchi K and Uryu K 2005 
Merger of binary neutron stars with realistic equations of state 
in full general relativity 
\emph{Phys. Rev. D} {\bf 71} 084021

\bibitem{villiers1}
De Villiers J-P, Hawley J F and Krolik J H 2003
Magnetically driven accretion flows in the Kerr metric 
I: Models and overall structure 
\emph{ApJ} {\bf 599} 1238

\bibitem{villiers2}
Hirose S, Krolik J H,  De Villiers J-P and Hawley J F 2004
Magnetically driven accretion flows in the Kerr metric 
II: Structure of the magnetic field
\emph{ApJ} {\bf 606} 1083

\bibitem{villiers3}
De Villiers J-P, Hawley J F, Krolik J H and Hirose S 2005
Magnetically driven accretion in the Kerr metric 
III: Unbound outflows
\emph{ApJ} {\bf 620} 878

\bibitem{mckinney1}
J C McKinney J C and Gammie C F 2004
A measurement of the electromagnetic luminosity of a Kerr black hole
\emph{ApJ} {\bf 611} 977

\bibitem{macfadyen1}
Proga D, MacFadyen A I, Armitage P J and Begelman M C 2003 
Axisymmetric MHD simulations of the collapsar model for gamma-ray bursts 
\emph{ApJ} {\bf 599} L5

\bibitem{macfadyen2}
Zhang W, Woosley S E and MacFadyen A I 2003 
Relativistic jets in collapsars 
\emph{ApJ} {\bf 586} 356

\bibitem{andersson1} 
Andersson N and Glampedakis K 2000
Superradiance resonance cavity outside rapidly rotating black holes 
\emph{Phys. Rev. Lett.} {\bf 84} 4537

\bibitem{andersson2}
Glampedakis K and Andersson N 2001 
Late-time dynamics of rapidly rotating black holes 
\emph{Phys. Rev. D} {\bf 64} 104021

\bibitem{superrad1}
Krivan W, Laguna P and Papadopoulos P 1996 
Dynamics of scalar fields in the background of rotating black holes 
\emph{Phys. Rev. D} {\bf 54} 4728

\bibitem{superrad2}
Andersson N, Laguna P and Papadopoulos P 1998
Dynamics of scalar fields in the background of rotating black holes. 
II. A note on superradiance 
\emph{Phys. Rev. D} {\bf 58} 087503

\bibitem{detweiler1}
Detweiler S 1980 
Klein-Gordon equation and rotating black holes 
\emph{Phys. Rev. D} {\bf 22} 2323

\bibitem{strafuss1}
Strafuss M J and Khanna G 2005 
Massive scalar field instability in Kerr spacetime 
\emph{Phys. Rev. D} {\bf 71} 024034

\bibitem{press1}
Press W H and Teukolsky S A 1972 
Floating orbits, superradiant scattering and the black-hole bomb 
\emph{Nature} {\bf 238} 211

\bibitem{scalar1}
Beck C 2004 
Chaotic scalar fields as models for dark energy 
\emph{Phys. Rev. D} {\bf 69} 123515

\bibitem{scalar2}
Beck C 2004 
Nonextensive scalar field theories and dark energy models 
\emph{Physica A} {\bf 340} 459

\bibitem{scalar3}
Baccigalupi C, Balbi A, Matarrese S, Perrotta F and Vittorio N 2002 
What's behind acoustic peaks in the Cosmic Microwave Background 
Anisotropies 
\emph{Nuc. Phys. B} (Proc. Suppl.) {\bf 110} 173

\bibitem{scalar4}
Guth A H 1981 
Inflationary universe: A possible solution to the horizon and 
flatness problems 
\emph{Phys. Rev. D} {\bf 23} 347

\bibitem{scalar5}
Albrecht A and Steinhardt P J 1982 
Cosmology for Grand Unified Theories with radiatively induced symmetry 
breaking
\emph{Phys. Rev. Lett.} {\bf 48} 1220

\bibitem{weinberg0}
Weinberg S 1972 
{\em Gravitation and Cosmology} 
(New York: John Wiley \& Sons)

\bibitem{simu}
Scheel M A, Erickcek A L, Burko L M, Kidder L E, Pfeiffer H P 
and Teukolsky S A 2004 
3D simulations of linearized scalar fields in Kerr spacetime 
\emph{Phys. Rev. D} {\bf 69} 104006

\bibitem{starobinskii1}
Starobinsky A A and Churilov S M 1974 
Amplification of electromagnetic and gravitational waves scattered by 
a rotating black hole 
\emph{Sov. Phys. JETP} {\bf 38} 1

\bibitem{starobinskii2}
Starobinsky A A 1973 
Amplification of waves during reflection from a rotating black hole 
\emph{Sov. Phys. JETP} {\bf 37} 28

\bibitem{zeldovich1}
Zeldovich Ya B 1972 
Amplification of cylindrical electromagnetic waves reflected 
from a rotating body 
\emph{Sov. Phys. JETP} {\bf 35} 1085

\bibitem{stuchlik1}
Stuchl\'{i}k Z, Bao G, Ostgaard E and Hled\'{i}k S 1998 
Kerr-Newman-de Sitter black holes with a restricted repulsive barrier of 
equatorial photon motion 
\emph{Phys. Rev. D} {\bf 58} 084003

\bibitem{stuchlik2}
Stuchl\'{i}k Z and Hled\'{i}k S 2000
Equatorial photon motion in the Kerr-Newman spacetimes with a non-zero 
cosmological constant 
\emph{Class. Quant. Grav.} {\bf 17} 4541

\bibitem{stuchlik3}
Stuchl\'{i}k Z and Slan\'{y} P 2004 
Equatorial circular orbits in the Kerr-de Sitter spacetimes 
\emph{Phys. Rev. D} {\bf 69} 064001

\bibitem{gh1}
Gibbons and Hawking S W 1977 
Cosmological event horizons, thermodynamics, and particle creation 
\emph{Phys. Rev. D} {\bf 15} 2738

\bibitem{gibbons1}
Gibbons G W, Lu H, Page D N and Pope C N 2005 
The general Kerr-de Sitter metrics in all dimensions
\emph{J. Geom. Phys.} {\bf 53} 49

\bibitem{vasudevan1}
Vasudevan M, Stevens K A and Page D N 2005 
Separability of the Hamilton-Jacobi and Klein-Gordon equations 
in Kerr-de Sitter metrics 
\emph{Class. Quant. Grav.} {\bf 22} 339

\bibitem{suzuki1}
Suzuki H, Takasugi E and Umetsu H 1999 
Analytic solutions of Teukolsky equation in Kerr-de Sitter and 
Kerr-Newman-de Sitter geometries
\emph{Prog. Theor. Phys.} {\bf 102} 253

\bibitem{suzuki2}
Suzuki H, Takasugi E and Umetsu H 1998 
Perturbations of Kerr-de Sitter black hole and Heun's equations
\emph{Prog. Theor. Phys.} {\bf 100} 491

\bibitem{fackerell}
Fackerell E D and Crossman R G 1977 
Spin-weighted angular spheroidal functions 
\emph{J. Math. Phys.} {\bf 18} 1849

\bibitem{seidel}
Seidel E 1989 
A comment on the eigenvalues of spin-weighted spheroidal functions 
\emph{Class. Quant. Grav.} {\bf 6} 1057

\bibitem{kraniotis1}
Kraniotis G V 2004 
Precise relativistic orbits in Kerr and Kerr-(anti)-de Sitter spacetimes 
\emph{Class. Quant. Grav.} {\bf 21} 4743

\bibitem{sandi2}
Setiawan S 
\emph{in preparation}

\bibitem{sandi}
Setiawan S, Ruffert M and Janka H-Th 2004
Non-stationary hyperaccretion of stellar-mass black holes in three dimensions: 
Torus evolution and neutrino emission 
\emph{MNRAS} {\bf 352} 753

\bibitem{rees}
Begelman M C, Blandford R D and Rees M J 1984 
Theory of extragalactic radio sources 
\emph{Rev. Mod. Phys.} {\bf 56} 255

\bibitem{reva1}
Williams R K 1995 
Extracting x rays, gamma rays, and relativistic e$^-$e$^+$ pairs from 
supermassive Kerr black holes using the Penrose mechanism
\emph{Phys. Rev. D} {\bf 51} 5387

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{fullpage}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage[dvips]{epsfig}

\def\##1{\underline{#1}}
\def\=#1{\underline{\underline{#1}}}

\def\+#1{\underline{\bf #1}}
\def\*#1{\underline{\underline{\bf #1}}}

\def\le{\left(}
\def\ri{\right)}
\def\les{\left[}
\def\ris{\right]}

\def\.{\mbox{ \tiny{$^\bullet$} }}

\begin{document}

\begin{center}

{\bf {\Large A short note on climbing gravity improbable :\\ Superradiant ring fellowship of launching\\ the high Lorentz-factor outflows/jets}}

 \vspace{4mm} \large

Sandi Setiawan\footnote{Fax: + 44 131 650 6553; e--mail:
S.Setiawan@ed.ac.uk}\\
{\em School of Mathematics,
University of Edinburgh, Edinburgh EH9 3JZ, UK}\\

\end{center}

\vspace{4mm}

\normalsize

\begin{abstract}
\noindent 
The spin energy extraction from a rotating black hole by amplification 
of scalar waves inside superradiance resonance wedge cavity (ring), 
followed by jet formation via magnetic reconnection assisted by 
the amplified scalar fields, is proposed. 
This mechanism may explain the availability of energy for as well as 
the formation of relativistic jets with high Lorentz-factor 
in microquasars and quasars. 
It is speculated that it may also explain 
that the spin of the central black hole, 
which facilitates the formation of superradiance cavity, 
which in turn would produce 
the amplified scalar waves by taking the hole's spin energy, 
is the source for the distinction between 
radio-loud and radio-quiet active galaxies. 

\end{abstract}

\noindent {\bf Keywords:} Superradiance, massless scalar fields, relativistic jets, gamma-ray bursts, radio active galaxies, black holes, Kerr-Newman-de Sitter spacetime

\vspace{4mm}

\section{Introduction}

Recently it was suggested that superradiant scattering of magnetosonic waves in a toroidal magnetosphere around a Kerr black hole may play a significant role in providing the energy for gamma-ray bursts \cite{putten1}. Using the WKB approximation, the amplification of scalar waves confined in a equatorial cavity wedge was computed and it was found that the cavity which acts as a waveguide for magnetosonic waves may give rise to instabilities in all superradiant modes. The author then suggested that the result for scalar waves should generalize to electromagnetic waves. However, this was not the case, though it was found that the scalar waves were amplified more strongly \cite{aguirre1} than the results obtained in \cite{putten1}.

It is also well-known that the black hole rotational energy may be extracted via Penrose process \cite{penrose1,bardeen1,wilkins1,wald1} (see also \cite{christodoulou1,christodoulou2,teukolsky0,teukolsky1,teukolsky2,teukolsky3,chandra0,aliev1}). Recent simulations showed that the so-called magnetohydrodynamical (MHD) Penrose process of extracting spin energy of a rotating black hole is only true for a short period of time when a large fraction of magnetic fields anchored in the black hole ergosphere \cite{komissarov1}, contrary to the claim suggested previously that this process may play a quite important role \cite{koide1}. The magnetic field lines are quickly pulled out from the ergosphere and moved into the black hole, and the MHD Penrose process almost completely stopped to operate \cite{komissarov1}. However, the extraction of rotational energy of the black hole is still taking place via Blandford-Znajek mechanism \cite{blandford1,znajek1} (see also \cite{blandford2,znajek2,znajek3}; and \cite{wilson1,xinli1,wald2,king1,petterson1,petterson2,chitre1} as well as \cite{bhmag1,bhmag1a,bhmag2,bhmag3,bhmag4}), which is believed to be responsible for the creation of jets (see, for example, \cite{narayan1,semenov1} and related references therein), that may be originated from within a few Schwarschild radius from the central black hole \cite{wilms1,iwasawa1,dabrowski1}. Unfortunately, from the results of advanced recent simulations (see, for example, \cite{komissarov1}) we learn that relativistic as well as collimated jets are failed to attain, and one should look for other explanations. However, there are still claims and counter-claims going on about the subject (see, for example, \cite{komissarov1,koide1,shibata1,villiers1,villiers2,villiers3,mckinney1,macfadyen1,macfadyen2}).

It was also discovered recently that a rapidly rotating black hole may have the so-called `superradiance resonance cavity' just outside it \cite{andersson1,andersson2} (see also \cite{superrad1,superrad2}). The existence of this cavity may physically explain the presence of large number of slowly damped quasi-normal modes (QNMs) trapped inside it \cite{andersson1}. If this kind of cavity happens to be physically similar to the wedge cavity in \cite{putten1}, in which it is proposed that it resembles the vacuum cavity around a black hole enclosed by a mirror as found in \cite{press1}, then it would be worth exploring the possibility of extracting spin energy from a rotating black hole via amplification of scalar waves trapped inside the cavity, which then followed by jet formation via magnetic reconnection assisted by those amplified scalar fields.

\section{Theoretical framework}

Massive scalar fields in Kerr spacetime background are unstable and their amplitudes can grow without bound \cite{detweiler1,strafuss1}. However, the instability growth time is too short to be astrophysically important \cite{detweiler1}. For the massless scalar fields, which are very important in the proposed scenario explored in this paper, one can find interesting discussion in \cite{press1}; scalar fields as dark energy which produce accelerating environment can be found in \cite{scalar1,scalar2,scalar3} (see \cite{scalar4,scalar5} for scalar fields in the context of inflationary universe model; see also \cite{weinberg0} for general account on scalar fields and \cite{simu} for a version of its simulations). Related to this topic, one can also find earlier accounts on this subject in \cite{starobinskii1,starobinskii2,zeldovich1}. 

We will not do detailed analysis in this paper. What will be done here is only to propose a theoretical framework on which we may build up a scenario about what physical mechanism that may be responsible for launching the high Lorentz-factor outflows/jets. First, we consider a rotating Kerr black hole. The Kerr black hole would transform itself into the Kerr-Newman hole if it is immersed in the magnetic fields since the induced charges would be acquired in this situation (see, for example, \cite{xinli1,wald2}). We then investigate the perturbation of massless scalar fields in this Kerr-Newman spacetime \cite{teukolsky0,teukolsky1,teukolsky2,teukolsky3} (see also \cite{fackerell,seidel}). According to the authors of \cite{andersson1,andersson2}, superradiance resonance cavity would be present in this situation and it is located just outside the black hole's outer event horizon (see also \cite{superrad1,superrad2}). We compare the result obtained from that of Kerr-Newman spacetime with the one from a very narrow equatorial wedge background \cite{putten1,aguirre1}. We would then consider a kind of magnetized Kerr-Newman-de Siter spacetime (see, for example, \cite{stuchlik1,stuchlik2,stuchlik3,gh1,gibbons1,vasudevan1,suzuki1,suzuki2}) as the relevant background due to the presence of the amplified massless scalar waves caused by the superradiance cavity to investigate the motion of the particles/waves/radiation in the vicinity of the black hole.

The Kerr-Newman-de Sitter metric in the Boyer-Lindquist coordinates \cite{gh1,suzuki1,suzuki2} can be expressed as follows : 

\begin{eqnarray}
\label{bl}
ds^2 &=& -\rho^2
\left(\frac{dr^2}{\Delta_r}+\frac{d\theta^2}{\Delta_\theta}\right)
-\frac{\Delta_\theta \sin^2\theta}{(1+\frac{\displaystyle \Lambda a^2}{\displaystyle 3})^2 \rho^2}
[a \, dt-(r^2+a^2) \, d\varphi]^2 \nonumber \\ 
&&  +\frac{\Delta_r}{(1+\frac{\displaystyle \Lambda a^2}{\displaystyle 3})^2 \rho^2}(dt-a\sin^2\theta \, d\varphi)^2,
\end{eqnarray}
with 
\begin{eqnarray}
\Delta_r &=& \le r^2+a^2 \ri \le 1-\frac{\Lambda r^2}{3} \ri -2Mr+ Q^2 \,, \\
\Delta_\theta &=& 1+\frac{\Lambda}{3} a^2 \cos^2\theta \,, \\
\rho^2 &=& r^2 + a^2 \cos^2\theta \,, 
\end{eqnarray}
wherein $\Lambda$ is the term (that behaves like `cosmological constant') which represents the amplified uncharged massless scalar fields, $M$ the mass of the black hole, $a$ its angular momentum per unit mass and $Q$ its charge. However, it is very unlikely that $\Lambda$ would be constant in this case since the scalar fields may be produced differently depending on their distance from the hole's outer event horizon (perhaps, for example, $\Lambda \propto \frac{\displaystyle 1}{\displaystyle r}$), and the scalar fields themselves would also evolve with time.

Since we are dealing with a Kerr rotating black hole immersed in magnetic fields, we would consider the Kerr-Newman background because the Kerr black hole would acquire the induced charge caused by the magnetic fields. We then investigate separated equations for the perturbation of uncharged massless fields with arbitrary spin $s$ in Kerr-Newman background using Boyer-Lindquist coordinates. We can simply take $\Lambda = 0$ in this case. And, we get the angular Teukolsky equation \cite{teukolsky0} (see also \cite{aguirre1}) :

\begin{equation}
\label{eqn:sx}
\les \frac{1}{\sin\theta}\frac{d}{d\theta}\sin\theta \frac{d}{d\theta} + a^2\omega^2\cos^2\theta + 2 a \omega s \cos\theta -\frac{(m+s\cos\theta)^2}{\sin^2\theta} + E-s^2  \ris S = 0 \,,
\end{equation}
and the radial Teukolsky equation is given by \cite{teukolsky0} (see also \cite{aguirre1}) :
\begin{equation}
\les \Delta^{-s}\frac{d}{dr}\Delta^{s+1}\frac{d}{dr} +
  \frac{K^2+2is(r-M)K}{\Delta} -4ir\omega s 
-E-2am\omega-a^2\omega^2+ s(s+1) \ris R = 0 \,, 
\end{equation}
wherein $\Delta \equiv r^2 - 2Mr + a^2 + Q^2$ and $K \equiv \le r^2+a^2 \ri \omega + am$, with $E$ an angular eigenvalue.

Chandrasekhar showed that the radial equation can be expressed as a wave equation with a potential barrier \cite{chandra0}. In this paper, we are interested in the uncharged massless scalar fields, i.e., $s=0$. For this case, the potential barrier is given by :

\begin{equation}
V = \frac{\Delta}{\rho^4} \les \lambda + \frac{1}{\rho^2} \le \Delta + 2r \le r-M \ri \ri -3 \frac{r^2\Delta}{\rho^4} \ris \,,
\end{equation}
and $\lambda$ is the angular eigenvalue given by $\lambda = E+a^2\omega^2+2am\omega$, in which $E$ is \cite{fackerell,seidel} (see also \cite{aguirre1}) :

\begin{equation}
E = l(l+1) + 2a^2\omega^2 \les \frac{m^2-l(l+1)+\frac{1}{2}}{(2l-1)(2l+3)} \ris + O\les (a\omega)^4 \ris \,,
\end{equation} 
which is obtained from the boundary conditions of regularity at $\theta=0$ and $\theta=\pi$ (see \cite{aguirre1}). This value should then be compared with the one obtained by assuming a constant angular function across the narrow superradiance resonance wedge cavity, which is 
\begin{equation}
E \simeq m^2 + s^2 \,,
\end{equation}
by taking $dS/d\theta=0$ in equation (\ref{eqn:sx}), followed by taking the approximation $\cos\theta \rightarrow 0$ and $\sin\theta \rightarrow 1$ (see \cite{putten1} and also \cite{aguirre1}). 

For the scalar case where $s=0$, 
the very narrow wedge geometry changes the angular eigenvalue for $m=1$
from $E = 2+O[(a\omega)^2]$ to $E=1$; the eigenvalue would add to the
height of the potential barrier, so the lower eigenvalue in the wedge
geometry would produce much stronger superradiance \cite{aguirre1}. 
We should also note that the presence of the charge $Q$ reduces the 
size of the outer event horizon and the static limit of the hole.

We should then be able to analyze the motion of the particles/waves/radiation 
in the vicinity of the black hole immersed in magnetic fields (which may be 
driven by the currents in the accretion disk surrounded the hole) 
using a more general magnetized Kerr-Newman-de Sitter spacetime 
(see, for example, \cite{bhmag1a}; see also \cite{kraniotis1}) with 
$\Lambda$ inconstant, in order to establish the right environment for 
launching a relativistic outflows/jets. Detailed accounts will be 
published elsewhere \cite{sandi2}.

We note that 
this superradiance effect would take the spin energy of the black hole and 
create the background of (amplified) scalar fields which is similar to 
the de Sitter spacetime (these scalar fields may also be interpreted as
something similar to dark energy in an accelerating de Sitter-like universe
-- see, for example, \cite{scalar1,scalar2,scalar3}). 
Combining this background with the spacetime created by 
the charged-induced rotating black hole \cite{xinli1} 
would result in a spacetime with a kind of 
Kerr-Newman-de Sitter metric especially in a region close to the black hole. 
The difficulties of having relativistic flows and jets are solved 
by the fact that it would be natural for matter and radiation near 
the black hole to attain high-Lorentz factor velocity by `riding' 
the amplified massless scalar waves created 
in the ergosphere of a rotating black hole
which provide the de Sitter-like spacetime environment which is 
accelerating. 
This process may explain the availability of energy for as well as 
the formation of relativistic jets with very high-Lorentz factor 
in microquasars 
(for example, in form of gamma-ray bursts \cite{sandi}) and quasars \cite{rees}. 

In addition, the generated magnetic reconnections (in relation with 
the matter of the accreted disk) may also help to collimate the jets. 
This mechanism may also explain that the spin of the central black hole, 
which facilitates the formation of superradiance cavity, 
which in turn would produce the amplified uncharged massless 
scalar fields by taking the hole's spin energy, 
is the source for the distinction between 
radio-loud and radio-quiet active galaxies \cite{wilson1}. 
It may also be responsible for the presence of the dark energy 
(or/and, perhaps, dark matter) in our universe.
We also note that the scenario proposed here is different from that of 
\cite{reva1}.

\section{Concluding remarks}

The proposal considered in this paper is still in its very early development. Further consideration, both analytic and numerical, needs to be explored in order to confirm the claims made in this paper. It is very tempting to know what large-scale numerical simulations would say about this by considering the full magnetohydrodynamical general relativistic approach and by also taking into account the effects arise from the presence and evolution of superradiance wedge cavity (ring) in the ergosphere of a rotating black hole.

\bigskip

\noindent{\bf Acknowledgement}

\noindent The author acknowledges EPSRC for support under grant GR/S60631/01.


\end{document}